# Beijing 100084, China

### v1

@@ IV-B Algorithm Derivation and Implementation
By introducing the multipliers $\bm{\alpha}\in\mathbb{R}^{M},\lambda\in\mathbb{R}^{+},\eta\in\mathbb{R}$, the Lagrangian of the semi-relaxed RDD function (7) is:
@@
Our key ingredient is optimizing the primal variables $W,\bm{r}$ in an alternative manner. Based on the Lagrangian of the RDD function, we take derivatives with respect to the primal variables $W$ and $\bm{r}$ for its optimal expression.
a) Updating $W=(w_{ij})$: When taking the derivative of $\mathcal{L}(W,\bm{r};\bm{\alpha},\lambda,\eta)$ with respect to the primal variable $W$, to address the first-order condition, we utilize the mirror descent method to linearize the quadratic component of $W$, one has
@@
Thus $w_{ij}^{(k+1)}$ can be expressed in the closed-form of
@@
To summarize, we update the variables $W$ and $\bm{r}$ in an alternating manner. For clarity, the pseudo-code is presented in the below Algorithm 1.
@@
0:  marginal distribution $P_{X}$, maximum iteration $max\_iter$, metric space $(X,d_{\mathcal{X}}),(Y,d_{\mathcal{Y}})$ and metric matrix $D^{\mathcal{X}},D^{\mathcal{Y}}$
@@
The computational complexity of the proposed algorithm is $\mathcal{O}(M^{2}N+MN^{2}))$. The primary bottleneck occurs in line 5, which involves matrix multiplication. In contrast, all other operations consist solely of vector multiplications.
@@ V Numerical Experiments and Discussions
This section investigates the relevant properties of the RDD function through numerical experiments. For starters, we compute the RDD function of the discrete versions of three classical distributions (i.e., the Gaussian source, Laplacian source and uniform source) across various metric spaces. All three sources utilize the square of $L_{2}$ norm distance, i.e., $q=2$. For simplicity, we refer to $dim(\mathcal{X})$ as the dimension of the metric space $\mathcal{X}$. For the implementation of the proposed AMD algorithm, the multiplier $\lambda$ is chosen as an arithmetic sequence of $100$ elements and we compute the result for each given $\lambda$. The parameter $max\_iter$ is set to $100$.
First, we consider the case where points in the metric measure spaces $\mathcal{X}$ and $\mathcal{Y}$ are distributed in a uniform grid. Specifically, we truncate the sources of each dimension into an interval $[-h,h]$ and then discretize the interval by a set of uniform grid points: $x_{i}=-h+\delta/2+(i-1)\cdot\delta$, $\delta=2\cdot h/K$, $i=1,\cdots,K,$ where $K$ denotes the number of coordinates of each dimension. Therefore, the two-dimensional and three-dimensional metric spaces contain $K^{2}$ and $K^{3}$ points, respectively. The probability distribution $P_{X}$ is generated from the density functions of sources and then normalized. In the sequel, we take $h=8$ for all three sources.
Figure 1: The RDD functions of different sources and different metric space dimensions. Left: $dim(\mathcal{X})=dim(\mathcal{Y})=1$; Right: $dim(\mathcal{X})=2,dim(\mathcal{Y})=3$.
It can be observed that the proposed AMD algorithm demonstrates the convergence property to some extent for discrete sources.
Next, we simulate the point distributions on spherical surfaces i.e., a circle and a sphere, each with a radius of $4$ units. For the circle, we calculate the angular increments to evenly distribute $20$ points around its perimeter and use trigonometric functions to obtain their Cartesian coordinates. As for the sphere, we establish the angular positions for $20$ points across each latitude and longitude, employing a meshgrid to create a grid that covers the sphere’s surface comprehensively. This approach yields a total of $20$ points on a circle and $400$ points on a sphere, respectively. The corresponding RDD curve is plotted in Fig. 2. This illustration verifies that the proposed RDD function and AMD algorithm extend their applicability beyond merely uniform grids.
Finally, we discuss the relationship between RDD and RD functions by considering the following problem, i.e., the constraints involve both distortion and distortion-in-distortion,
@@
$+(1-\theta)\mathbb{E}[d(X,Y)]\leq D,(X,Y,X^{\prime},Y^{\prime})\sim P_{XY}\times P_{XY},$ (8b)
Figure 2: The RDD function of Gaussian source on spherical surfaces. The dimensions of metric spaces are $dim(\mathcal{X})=2,dim(\mathcal{Y})=3$.
The constraint (8b) draws inspiration from the form of Fused GW [21], which combines Wasserstein and Gromov-Wasserstein distances to jointly account for the features and structures of graphs. The function (8) can be solved by slight modifications to the proposed AMD method introduced in Section IV while incorporating the idea of the Blahut-Arimoto (BA) algorithm [22]. The detailed algorithmic procedure is provided in the Appendix . Here, we consider a one-dimensional Gaussian distribution with $\sigma=2$ and set the distortion measure $d$ as the squared error.
Figure 3: The $R(D;\theta)$ surface and its corresponding cross-section curves. Left: $R(D;\theta)$ surface with $\theta\in[0,0.1]$; Middle: $R$-$D$ curve; Right: $R$-$\theta$ curve. The dimensions of metric spaces are $dim(\mathcal{X})=dim(\mathcal{Y})=1$.
Fig. 3 illustrates the $R(D;\theta)$ surface along with its corresponding cross-section curves. Whether we examine the overall surface (left), the $R$-$D$ curve with a fixed $\theta$ (middle), or the $R$-$\theta$ curve with a fixed D (right), it is apparent that as $\theta$ increases (indicating a growing influence of distortion-in-distortion) and $D$ decreases, the $R$-value increases monotonically. This indicates that the impact of the distortion-in-distortion term on the distortion term is significant. We acknowledge that the distortion-in-distortion term lacks theoretical convexity. However, even when $\theta$ is very small (middle), where the convex distortion term dominates, the results of numerical experiments clearly demonstrate the monotonic relationship of $R$ concerning both $\theta$ and $D$. Due to space limitations, the conditional probabilities $P_{Y|X}$ for various values of $\theta$ and $D$ are presented in Appendix . Additionally, we can observe how changes in $\theta$—specifically, changes in the weight of the distortion-in-distortion term—affect the conditional probability $P_{Y|X}$.
@@ VI Conclusion
This paper introduces a novel function i.e., RDD function, which serves as an extension of the classical RD function. The theoretical validity of the RDD function is established through a coding theorem, demonstrating its rationale as an operational rate-distortion function. In order to conquer the high computational complexity caused by the Gromov-type distortion and to efficiently compute the RDD function, we propose a low-complexity AMD algorithm and test it on several classical sources for various metric spaces. This work primarily presents a theoretical exploration, while its practical applications remain under investigation. As an extension of the classical RD function—just like the RDP function—we believe that the RDD function also holds promise for applications in a wide range of fields, such as the point cloud compression.
@@
0:  marginal distribution $P_{X}$, maximum iteration $max\_iter$, metric space $(X,d_{\mathcal{X}}),(Y,d_{\mathcal{Y}})$ and metric matrix $D^{\mathcal{X}},D^{\mathcal{Y}}$

### v2

@@ IV-B Algorithm Derivation and Implementation
By introducing multipliers $\bm{\alpha}\in\mathbb{R}^{M},\lambda\in\mathbb{R}^{+},\eta\in\mathbb{R}$, the Lagrangian of the semi-relaxed RDD function (8) is:
@@
Our key idea is optimizing the primal variables $W,\bm{r}$ in an alternative manner. Based on the Lagrangian of the RDD function, we take derivatives with respect to the primal variables $W$ and $\bm{r}$ for their optimal expression.
a) Updating $W=(w_{ij})$: When taking the derivative of $\mathcal{L}(W,\bm{r};\bm{\alpha},\lambda,\eta)$ with respect to the primal variable $W$, to address the first-order condition, we utilize the mirror descent method to linearize the quadratic component of $W$, to get
@@
Thus $w_{ij}^{(k+1)}$ can be expressed in closed form as
@@
To summarize, we update the variables $W$ and $\bm{r}$ in an alternating manner. For clarity, the pseudo-code is presented in Algorithm 1. Note that the multiplier $\lambda$ remains fixed, consistent with classical approaches like the Blahut-Arimoto (BA) algorithm [21].
@@
0:  marginal distribution $P_{X}$, maximum iteration $max\_iter$, metric spaces $(X,d_{\mathcal{X}}),(Y,d_{\mathcal{Y}})$ and metric matrices $D^{\mathcal{X}},D^{\mathcal{Y}}$
@@
The computational complexity of each iteration in Algorithm 1 is $\mathcal{O}(M^{2}N+MN^{2})$. The primary bottleneck occurs in line 5, which involves matrix multiplication. In contrast, all other operations consist solely of vector multiplications.
@@ V Numerical Experiments and Discussions
This section investigates the behaviors of the RDD function through numerical experiments. For starters, we compute the RDD function of the discrete versions of three classical sources (i.e., Gaussian, Laplacian and uniform) across various metric spaces. All three sources utilize the square of $L_{2}$ norm distance, i.e., $q=2$. For simplicity, we refer to $\rm{dim}(\mathcal{X})$ as the dimension of the metric space $\mathcal{X}$. For the implementation of the proposed AMD algorithm, the multiplier $\lambda$ is chosen as an arithmetic sequence of $100$ elements and we compute the result for each given $\lambda$. The parameter $max\_iter$ is set to $100$.
First, we consider the case where points on the metric measure spaces $\mathcal{X}$ and $\mathcal{Y}$ are distributed on a uniform grid. Specifically, we truncate the sources of each dimension into an interval $[-h,h]$ and then discretize the interval by a set of uniform grid points: $x_{i}=-h+\delta/2+(i-1)\cdot\delta$, $\delta=2\cdot h/K$, $i=1,\cdots,K,$ where $K$ denotes the number of coordinates of each dimension. Therefore, the two-dimensional and three-dimensional metric spaces contain $K^{2}$ and $K^{3}$ points, respectively. The probability distribution $P_{X}$ is generated from the density functions of sources and then normalized. In the sequel, we take $h=8$ for all three sources.
Figure 1: The RDD functions of different sources and different metric space dimensions. Left: $\rm{dim}(\mathcal{X})=\rm{dim}(\mathcal{Y})=1$; Right: $\rm{dim}(\mathcal{X})=2,\rm{dim}(\mathcal{Y})=3$. The parameter $\sigma$ denotes the scale parameter for both Gaussian and Laplacian sources, see Appendix for the definition of these sources due to space limitation.
The corresponding RDD curves of different sources and different metric space dimensions are presented in Fig. 1. As shown, the RDD curves obtained from the AMD algorithm demonstrate a decreasing trend similar to the classical RD curve.
Next, we consider discrete distributions on spherical surfaces i.e., circle (Fig. 2 Left) and sphere (Fig. 2 Right), each with a radius of $4$ units. For the circle, we calculate the angular increments to evenly distribute $20$ points around its perimeter and use trigonometric functions to obtain their Cartesian coordinates. As for the sphere, we establish the angular positions for $20$ points across each latitude and longitude, employing a meshgrid to create a grid that covers the sphere’s surface comprehensively. This approach yields a total of $20$ points on a circle and $400$ points on a sphere, respectively. The corresponding RDD curve is plotted in Fig. 2 Middle. This illustration verifies that the proposed RDD function and AMD algorithm extend their applicability beyond uniform grids.
Finally, we discuss the relationship between RDD and RD functions by considering the following problem, with the constraints involving both distortion and distortion-in-distortion.
@@
$+(1-\theta)\mathbb{E}[d(X,Y)]\leq D,(X,Y,X^{\prime},Y^{\prime})\sim P_{XY}\times P_{XY}.$ (9b)
Figure 2: The RDD function of Gaussian source on spherical surfaces and corresponding spatial configurations. Left: points on circle with dimension $\rm{dim}(\mathcal{X})=2$; Middle: $R$-$D$ curve; Right: points on sphere with dimension $\rm{dim}(\mathcal{Y})=3$.
The constraint (9b) draws inspiration from the form of Fused GW [22], which combines Wasserstein and Gromov-Wasserstein distances to jointly account for the features and structures of graphs. The function (9) can be solved by slight modifications to the proposed AMD method introduced in Section IV while incorporating the idea of the Blahut-Arimoto (BA) algorithm [21]. The detailed algorithmic procedure is provided in the Appendix. As an example, we consider a one-dimensional Gaussian distribution with $\sigma=2$ and set the distortion measure $d$ as the squared error.
Figure 3: The $R(D;\theta)$ surface and its corresponding cross-section curves. Left: $R(D;\theta)$ surface with $\theta\in[0,0.1]$; Middle: $R$-$D$ curve; Right: $R$-$\theta$ curve. The dimensions of metric spaces are $\rm{dim}(\mathcal{X})=\rm{dim}(\mathcal{Y})=1$.
Fig. 3 illustrates the $R(D;\theta)$ surface along with its corresponding cross-section curves. Whether we examine the overall surface (left), the $R$-$D$ curve with a fixed $\theta$ (middle), or the $R$-$\theta$ curve with a fixed $D$ (right), it is apparent that as $\theta$ increases (indicating a growing influence of distortion-in-distortion) and $D$ decreases, the $R$-value increases monotonically. This indicates that the impact of the distortion-in-distortion term on the distortion term becomes significant.
It is still unclear whether the RDD function is convex with respect to the RDD term. However, even when $\theta$ is very small (middle), where the convex distortion term dominates, the numerical results clearly demonstrate the monotonic relationship of $R$ concerning both $\theta$ and $D$. Due to space limitations, the conditional probabilities $P_{Y|X}$ for various values of $\theta$ and $D$ are presented in Appendix. Additionally, we can observe how changes in $\theta$—specifically, changes in the weight of the distortion-in-distortion term—affect the conditional probability $P_{Y|X}$.
@@ VI Conclusion
This paper proposes a novel function—the Rate Distortion-in-Distortion (RDD) function, extending the classical RD framework. Specifically, it replaces the conventional distortion constraint with Gromov-type distortion (which can be viewed as a structural distortion measure). The theoretical validity of the RDD function is established through a coding theorem, demonstrating its rationale as an operational information-theoretic measure. In order to tackle the high computational complexity caused by the Gromov-type distortion and to efficiently compute the RDD function, we propose a low-complexity AMD algorithm and test it on several classical sources for various metric spaces. In this preliminary stage, this work primarily presents a theoretical exploration, while its practical applications remain under investigation. As an extension of the classical RD function—just like the RDP function—we expect that the RDD function also holds promise for applications in a wide range of fields, such as point cloud compression, radio map representation, and generative models.
@@
0:  marginal distribution $P_{X}$, maximum iteration $max\_iter$, metric spaces $(X,d_{\mathcal{X}}),(Y,d_{\mathcal{Y}})$ and metric matrices $D^{\mathcal{X}},D^{\mathcal{Y}}$